Service Status Icon

Service Status

Check our service status and past incidents

Overall Service Status

100%

Up

All services are operating normally at this time.

Last Incident


Status: Resolved - Posted at: 10th August 2021 - 16:02

All Regions

Data Center Current Last Day Last Month Last Year
Dallas DC Icon
Dallas, TX
Up 100% 100% 100%
Seattle DC Icon
Seattle, WA
Up 100% 100% 100%
New Jersey DC Icon
Piscataway, NJ
Up 100% 100% 100%
Los Angeles DC Icon
Los Angeles, CA
Up 100% 100% 100%
London Datacenter Icon
London, UK
Up 100% 100% 100%
Maidenhead Datacenter Icon
Maidenhead, UK
Up 100% 100% 100%
Netherlands Datacenter Icon
The Netherlands
Up 100% 100% 100%

Past Incidents

Resolved

Packet Loss in Los Angeles

Posted at: 17th August 2014 - 11:58

We are experiencing some packet loss in this network. We are in contact with the DC. Please bare with us.
Resolved

Los Angeles High Storage Network is Down

Posted at: 15th August 2014 - 19:59

There is a network issue with our IP subnets in this network zone. We are in contact with the DC.

Update: It looks like we are getting affected by this issue: http://blogs.cisco.com/sp/global-internet-routing-table-reaches-512k-milestone/ and getting filtered by some of the internet carriers.

Update 23:08: The issue is temporarily fixed. You may experience slowness until tomorrow.
Resolved

Los Angeles Maintenance

Posted at: 11th August 2014 - 15:55

Date: 11 August 2014, Monday
Start Time: 2:00 PM PST
End Time: 2:30 PM PST
Cause: Expansion
Expected Downtime: 30 Minutes

In order for us to deploy new nodes in our Los Angeles network, we are going to have to move a few servers off their current rack to a new rack.

We apologize for the inconvenience this maintenance will cause.

Update 1:55 PM (PST): Maintenance has started, we are shutting down the VPS's.

Update 2:20 PM (PST): All VPS's are powered down, moving started.

Update 3:05 PM (PST): All the VPS's are back online.

Resolved

node2 Dallas is Down

Posted at: 7th August 2014 - 19:56

node2 Dallas just kernel panic'ed, we are rebooting it.

Update 20:10: All containers have booted. They'll go down for a few moments and come back online while the server recalculates the quotas used.

Update 20:40: Everything is back to normal.
Resolved

Los Angeles is down

Posted at: 18th July 2014 - 05:27

We are aware that Los Angeles nodes are down. There is a network issue which the DC is working on.

Update 05:30: The network is back up.
Resolved

Offloaded SQL server Maintenance

Posted at: 29th June 2014 - 10:03

From the announcement we sent out last week:

You are receiving this email because you have active offloaded MySQL service.

We are writing to inform you of a maintenance window regarding our Dallas offloaded MySQL service. This window is scheduled for Sunday, June 29th from 10 AM to 11 AM CDT (-5 UTC).

We do not expect to have a total outage of the service, but connectivity may be intermittent as we restart the SQL server process.

Please open a ticket if you have any questions.

Resolved

node2 issues

Posted at: 24th June 2014 - 17:24

Rebooting caused a kernel panic on this node2 Dallas and we are running fsck at the moment. Please be patient.

Update 19:16: fsck is done. The node is booted. VPSes are booting at the moment. To not to wait, go to your client area and boot your VPS.
Resolved

lapure1 does not come back online

Posted at: 24th June 2014 - 16:42

lapure1 is not coming back online. We are working on it. Your patience is appreciated.

Update: This node is back online.
Resolved

Emergency Node Reboots Due to Critical OpenVZ Vulnerability

Posted at: 24th June 2014 - 15:24

We are writing to inform you that we are going to reboot all our nodes in the following 15 minutes.

This is due to a critical OpenVZ vulnerability which is being discussed here: http://www.webhostingtalk.com/showthread.php?p=9155713

Please prepare for downtime.

Resolved

dalstorage1 is down

Posted at: 24th June 2014 - 08:02

dalstorage1 is facing network issues. We believe the DC's router is died. We are in contact with them.

Update 08:08 AM: All services are restored.
Resolved

node3 is down

Posted at: 22nd June 2014 - 18:06

node3 (Los Angeles) has just gone to kernel panic. We are investigating and restarting soon.

Update 18:11: Node has been restarted. The reason for kernel panic looks like a NIC driver bug. The VPSes are starting at the moment. This will take a few. Alternatively, you can start them on your client area.

Update 18:42: All VPSes are online.
Resolved

Urgent Network Maintenance in Dallas

Posted at: 19th June 2014 - 10:20

Our datacenter has notified us that there will be an urgent network maintenance today (06/19/2014) at 10 PM CST. They except a downtime of around 10 minutes.

This will affect all the High Memory plans in Dallas, TX location.

We will update this announcement once the maintenance is done.

We are sorry for the inconvenience this may cause.

Update: 10:20 PM: Network is back online.
Resolved

Los Angeles Network Issues

Posted at: 12th June 2014 - 07:46

There are routing issues within Asia to Los Angeles and affecting the connectivity. Note that, this is not something within our power. It's a carrier degree outage. We have notified the datacenter and asked them to update routes. However, we are not sure if they are going to do that for us.

You can see that your server is not actually offline by pinging its IP here: http://cloudmonitor.ca.com/en/ping.php

U
pdate 08:29 AM: Reply from the DC: "There appears that there was an issue with Coresite which has been resolved. I apologize for the inconvenience."
Resolved

LA node3 outage

Posted at: 10th June 2014 - 13:08

We are seeing an outage of our node3 server in LA. We have opened a ticket with the datacenter.
Resolved

ukpure1 NIC replacement

Posted at: 7th June 2014 - 22:38

We are taking the node down for a NIC (network) card replacement. It should be online in the following 15 mins.

Update 23:12: All VPS are up and running.
Resolved

dalstorage1 kernel panic

Posted at: 7th June 2014 - 15:09

dalstorage1 kernel panic'd. We have issued a reboot and waiting for it to boot.

Update 15:14: It's online now, containers are booting.
Resolved

node3 reboot

Posted at: 6th June 2014 - 10:31

As of 06/06/2014 10:31 CDT, we are rebooting node3. You will experience a loss of connectivity until this server returns.
Resolved

lastorage1 and lapure1 are down

Posted at: 31st May 2014 - 10:53

lastorage1 and lapure1 are down because of a network issue. We have contacted the datacenter and are waiting for a reply.

Update 11:46 AM: It seems on of the core routers failed with our datacenter and it's being replaced now.

Update 13:40 AM: Services are restored.
Resolved

ukpure1 has hardware problems

Posted at: 25th May 2014 - 22:20

There seems to be an issue with ukpure1 node. We believe it's a hardware problem. We are investigating.

Update 23:37: We believe we have fixed the issue and node seems to be stable at the moment.

Update May 27th 19:16: Issues came back. We are experiencing hardware problems at the moment.

Update 21:09: Issues are fixed.
Resolved

ukpure1 network issues

Posted at: 25th May 2014 - 21:11

ukpure1 is currently inaccessible due to network issues.

Edit: 3:40AM GMT+0 5/26/2014 resolved.

Contact Technical Support: